OpenStudy (anonymous):

A solution containing 45% water is mixed with another solution containing 10% bromine. If the chemist needs 30 liters of the mixture, how many liters of each solution are needed to make a 34.5% water/bromine solution?

OpenStudy (phi):

if we assume a bromine solution with 45% water, then it is a solution with 55% bromine call the amount of 55% bromide solution x call the amount of 10% bromide solution y when added together we want 30 L: x+y=30 we can figure out how much bromide we have: 0.55*x + 0.10*y = 0.345*30 (this assumes " 34.5% water/bromine solution" means 34.5% bromide...it is not clear what they mean) now you have 2 equations and 2 unknowns to solve.

OpenStudy (anonymous):

my final answer is 16 1/13 and 13 2/3

OpenStudy (phi):

I assume you mean 16 and 1/3 not 1/13 but yes, they look good.

OpenStudy (anonymous):

:)